Good day! can you please help me with this problem? thank you!
Find the quadratic equation with integral coefficients whose roots are twice the roots of: x^2 - 11x + 28 = 0

Answer by sachi(548)

x^2 - 11x + 28 = 0
or x^2-4x-7x+28=0
or x(x-4)-7(x-4)=0
(x-4)(x-7)=0
or x=4 or 7
the new roots=8,14
the quadratic eqn is
(x-8)(x-14)=0
or x^2-22x+112=0 ans